# OmniScience: A Domain-Specialized LLM for Scientific Reasoning and Discovery

### v2

@@ 3.1 Domain Adaptive Pretraining
Pretraining Data. Our pretraining corpus consists of 35 billion tokens collected from a diverse array of scientific sources, including 3.6 million peer-reviewed articles, as well as preprints from arXiv (Clement et al., 2019), ChemRxiv (Kiessling et al., 2016), PubChem (Kim et al., 2016), Semantic Scholar (Lo et al., 2020) and other such open research platforms. Additionally, the dataset integrates full text from academic textbooks. By combining natural language data from research papers and textbooks with structured scientific information, we capture a broad and comprehensive spectrum of scientific knowledge. This diversified data collection approach ensures a robust and representative foundation for modeling and analysis across scientific domains. Additional details about the domain adaptive pretraining data are provided in Sec. A.2 of the Appendix.
@@ 4.2 Results on Domain-Specific Benchmarks
We further evaluate our model using an SFT held-out test set (as detailed in Section 3.2) and an internally curated battery-specific reasoning benchmark. To ensure comprehensive comparison, we extend our analysis to large reasoning models with publicly accessible APIs, including GPT-o1, LLaMA 3.1 70B, Claude 3.7 Sonnet, and Gemini. As shown in Table 2, our OmniScience Reasoning model achieves robust performance across tasks such as Battery Q/A, multiple-choice questions (MCQ), reading comprehension, summarization, and reasoning, outperforming all baseline models except GPT-o1. Notably, while GPT-o1 (used to generate task instruction data for supervised fine-tuning) remains the strongest performer, our model achieves competitive results despite its significantly smaller parameter size and lower computational training costs. These findings highlight the efficacy of domain-adaptive pretraining and reasoning-focused distillation for specialized task performance.
@@ A.2 Domain Adaptive Data Collection
Peer-reviewed Papers 3.6M

### v4

@@ 3.1 Domain Adaptive Pretraining
Pretraining Data. Our pretraining corpus consists of 35 billion tokens collected from a diverse array of scientific sources, including large number of peer-reviewed articles, as well as preprints from arXiv (Clement et al., 2019), ChemRxiv (Kiessling et al., 2016), PubChem (Kim et al., 2016), Semantic Scholar (Lo et al., 2020) and other such open research platforms. Additionally, the dataset integrates full text from academic textbooks. By combining natural language data from research papers and textbooks with structured scientific information, we capture a broad and comprehensive spectrum of scientific knowledge. This diversified data collection approach ensures a robust and representative foundation for modeling and analysis across scientific domains. Additional details about the domain adaptive pretraining data are provided in Sec. A.2 of the Appendix.
@@ 4.1 Results on Public Benchmarks
Model MMLU Wino Hella ARC-E
LLaMA 3.1 70B 0.82 0.70 0.82 0.89
Omni-Chat 0.84 0.77 0.83 0.91
Omni-Reasoning 0.87 0.90 0.88 0.93
GPT-o1 0.91 0.95 0.95
DeepSeek-R1 0.91 0.95 0.98
Claude 3.7 (think) 0.86 0.89 0.95 0.96
Gemini 2.0 flash (think) 0.90 0.84 0.92 0.88
Grok 3 (think) 0.91
Table 2: Performance of SOTA reasoning models on four core public benchmarks — MMLU, Winogrande, Hellaswag, and ARC-E evaluating expert knowledge, commonsense, and reasoning. OmniScience Reasoning performs on par with the strongest models, which also demonstrates its robust generalization capabilities beyond the scientific domain.
Figure 3: Bar chart visualizing the performance of various LLMs on MMLU, Winogrande, Hellaswag, and ARC-E. OmniScience Reasoning consistently matches or exceeds stronger proprietary models across benchmarks.
In addition to GPQA Diamond, we benchmark OmniScience Reasoning on four widely used public benchmarks that assess general language understanding and reasoning, namely MMLU (Hendrycks et al., 2020), Winogrande (Sakaguchi et al., 2021), Hellaswag (Zellers et al., 2019), and ARC-E (Clark et al., 2018). We include these public benchmarks to demonstrate that OmniScience, while optimized for scientific reasoning, retains strong generalization capabilities. Table 2 summarizes the comparative performance on these benchmarks, and Figure 3 visualizes these results. Both the OmniScience Chat and Reasoning models consistently outperform the LLaMA-3.1-70B baseline, and OmniScience Reasoning remains competitive with significantly larger proprietary models, highlighting the effectiveness of our domain-adaptive and reasoning-based alignment strategies.
@@ 4.2 Results on Domain-Specific Benchmarks
We further evaluate our model using an SFT held-out test set (as detailed in Section 3.2) and an internally curated battery-specific reasoning benchmark. To ensure comprehensive comparison, we extend our analysis to large reasoning models with publicly accessible APIs, including GPT-o1, LLaMA 3.1 70B, Claude 3.7 Sonnet, and Gemini. As shown in Table 3 and Figure 4, our OmniScience Reasoning model achieves robust performance across tasks such as Battery Q/A, multiple-choice questions (MCQ), reading comprehension, summarization, and reasoning, outperforming all baseline models except GPT-o1. Notably, while GPT-o1 (used to generate task instruction data for supervised fine-tuning) remains the strongest performer, our model achieves competitive results despite its significantly smaller parameter size and lower computational training costs. These findings highlight the efficacy of domain-adaptive pretraining and reasoning-focused distillation for specialized task performance.
@@
Figure 4: Bar chart visualization of battery-specific task performance for various LLMs. This figure corresponds to Table 3 and highlights comparative accuracy across Q/A, MCQ, reading comprehension, summarization, and reasoning.
@@ A.2 Domain Adaptive Data Collection
Peer-reviewed Papers Few millions